# An Event Structure-aware Generative Model for Biomedical Event Extraction

### v2

@@ V-A Ablation Study
We conduct an ablation study of the proposed GenBEE model to evaluate the effects of different components on the overall performance across the MLEE, GE11, and PHEE benchmark datasets. Specifically, we focus on the following two components: Structural Prefix Learning and Event Prompt Construction. Table V reports the performance results of the ablation study. The symbol $\triangle$ indicates the difference in the F1-score between the different configuration models and the proposed GenBEE model. As shown in Table V, we observe that the structural prefixes and event prompts significantly contribute to the performance of our proposed model, regardless of the dataset. Firstly, the removal of the Structural Prefix Learning module from GenBEE results in a reduction of 1.5% and 1.6% in the average Trg-C and Arg-C F1-scores, respectively, highlighting the important role of this component in the model’s overall performance. Moreover, removing the event prompts leads to a decrease of 0.6% and 1.0% in the average Trg-C and Arg-C F1-scores, respectively. This drop in performance indicates that the prompts constructed for describing events significantly contribute to the improvement of the proposed GenBEE model. Overall, each component of the GenBEE model plays a crucial role in achieving promising performance for biomedical event extraction.
@@
(1) -w/o Structural Prefix Learning 79.8 66.7 66.4 62.1 68.7 52.6 -1.5 -1.6
(2) -w/o Event Prompts Construction 80.9 67.3 67.5 63.0 69.2 52.9 -0.6 -1.0
(3) -w/o (1) (2) 78.0 64.6 65.2 60.5 67.6 51.4 -2.9 -3.2
@@ V-B Performance Comparison between LLMs API and Fine-tuned PLMs
Given the immense potential of using large language models (LLMs) API with In-Context Learning (ICL) [21] across various NLP tasks under data efficient scenarios [38], we conduct experiments to compare the few-shot performance of the proposed GenBEE model with LLMs API. Specifically, we consider two widely used LLMs: GPT-4 [39] and Llama3-70b [40] for comparison. We access these LLMs through APIs from their official providers. As a part of the prompt, we provide LLMs with the same type of instructions, EE templates that we use in GenBEE, and a few ICL examples (positive ones). It is worth noting that the number of ICL examples is limited by the maximum context length supported by LLMs.
Table VI presents the performance results for LLMs and our proposed GenBEE model, categorized by shot numbers: 4-shot, 8-shot, 16-shot, and 32-shot. In the 4-shot learning scenario, the GPT-4 API achieves the highest F1 scores across all datasets, followed by the Llama3-70b API, with GenBEE showing the lowest performance. Similar trends are observed in the 8-shot and 16-shot learning scenarios, where all models exhibit performance improvements with increased learning data (shot numbers). However, for LLMs, adding more ICL examples beyond 16-shot does not necessarily result in further performance gains, as improvements tend to plateau or even slightly decline at higher shot levels. Notably, at 32-shot learning, GenBEE achieves the highest F1-scores across all datasets, particularly on the PHEE dataset with F1-scores of 43.7% and 38.5% for Trigger and Argument Classification, respectively. We infer that this is likely due to PHEE’s limited number of event types (only two), which reduces the number of unseen labels during few-shot training and allows GenBEE to leverage the provided examples more effectively. In comparison, datasets such as MLEE and GE11, with more diverse event types, present a greater challenge for the model, resulting in relatively lower performance gains. Generally, the fine-tuned GenBEE is able to outperform LLMs (using In-Context Learning) as the number of labeled data increases.
TABLE VI: Experimental results using LLMs API and GenBEE for few-shot learning based on the MLEE, GE11 and PHEE datasets. All reported numbers are in F1-scores (%).
@@ V-C Case Study
In this section, we compare GenBEE and DEGREE on two examples, which are presented in Figure 4, to illustrate the differences in their predictions. Example 1 presents a case where the DEGREE model incorrectly predicts the argument “Site2” in the Binding event triggered by “binding”, but our GenBEE model gives the correct prediction. We infer that our model has incorporated dependency information between arguments using event prompts for enhancing the performance of biomedical event extraction. Example 2 presents a case where the DEGREE model fails to predict the arguments “Theme” and “Site” in the Phosphorylation event, which is nested with the Binding event. However, the GenBEE model is able to predict the arguments of these two nested events correctly. We infer that our model has incorporated structural information using structural prefixes for effectively recognizing the relationships between nested events. As illustrated by the case study, with event prompts and structural prefixes, our proposed GenBEE model is able to perform biomedical event extraction effectively.
@@ VI Conclusion
We propose GenBEE, a novel generative model with structure-aware prefixes for biomedical event extraction. The experimental results demonstrate that GenBEE outperforms the strong baselines on the MLEE, GE11, and PHEE datasets. Moreover, our experiments show that structure-aware prefixes can effectively serve as a medium to link structural prompts with the representation space of generative models, thereby enhancing the model’s overall performance.

### v3

@@ V-A Ablation Study
We conduct an ablation study of the proposed GenBEE model to evaluate the effects of different modules on the overall performance across the MLEE, GE11, and PHEE benchmark datasets. Specifically, we focus on the following two modules: Event Prompt Construction and Structural Prefix Learning. Table V reports the performance results of the ablation study. The symbol $\triangle$ indicates the difference in the F1-score between the different configuration models and the proposed GenBEE model. As shown in Table V, we observe that event prompts and structural prefixes significantly contribute to the performance of our proposed model, regardless of the dataset. Removing the Event Prompt Construction module leads to a decrease of -0.6% and -1.0% on the average Trg-C and Arg-C F1-scores, respectively. This drop in performance indicates that the prompts constructed for describing events significantly contribute to the improvement of the proposed GenBEE model. Moreover, the removal of the Structural Prefix Learning module from GenBEE reduces the performance by -1.5% and -1.6% on the average Trg-C and Arg-C F1-scores, respectively, highlighting the important role of this module in the model’s overall performance. Finally, removing both Event Prompt Construction and Structural Prefix Learning modules will further degrade the model’s performance. Overall, each module of the GenBEE model plays a crucial role in achieving promising performance for biomedical event extraction.
@@
(1) w/o Event Prompt Construction 80.9 67.3 67.5 63.0 69.2 52.9 -0.6 -1.0
(2) w/o Structural Prefix Learning 79.8 66.7 66.4 62.1 68.7 52.6 -1.5 -1.6
(3) w/o (1) and (2) 78.0 64.6 65.2 60.5 67.6 51.4 -2.9 -3.2
@@ V-B Performance Comparison between LLMs API and Fine-tuned PLMs
Given the immense potential of using large language models (LLMs) API with In-Context Learning (ICL) [21] across various NLP tasks under data efficient scenarios [38], we conduct experiments to compare the few-shot performance of the proposed GenBEE model with LLMs API. Specifically, we consider two widely used LLMs: GPT-4 [39] and Llama3-70b [40] for comparison. We access these LLMs through APIs from their official providers. As a part of the prompt, we provide LLMs with the same type of instructions, EE templates that we use in GenBEE, and a few ICL examples (positive ones). It is worth noting that the number of ICL examples is restricted by the maximum context length supported by LLMs.
Table VI presents the performance results for LLMs and our proposed GenBEE model, categorized by shot numbers: 4-shot, 8-shot, 16-shot, and 32-shot. In the 4-shot learning scenario, the GPT-4 API achieves the highest F1 scores across all datasets, followed by the Llama3-70b API, with GenBEE showing the lowest performance. Similar trends are observed in the 8-shot and 16-shot learning scenarios, where all models exhibit performance improvements with increased learning data (shot numbers). However, for LLMs, adding more ICL examples beyond 16-shot does not necessarily result in further performance gains, as improvements tend to plateau or even slightly decline at higher shot levels. Notably, at 32-shot learning, GenBEE achieves the highest F1-scores across all datasets, particularly on the PHEE dataset with F1-scores of 43.7% and 38.5% for Trigger and Argument Classification, respectively. We infer that this is likely due to PHEE’s limited number of event types (only two), which reduces the need for more number of unseen labels during few-shot training and allows GenBEE to leverage the provided examples more effectively. In comparison, datasets such as MLEE and GE11 which contain more diverse event types present a greater challenge for the GenBEE model. Generally, the fine-tuned GenBEE with few-shot learning is able to outperform LLMs (using In-Context Learning) when the number of labeled data increases.
TABLE VI: Experimental results using LLMs API and GenBEE with few-shot learning based on the MLEE, GE11 and PHEE datasets. All reported results are in F1-scores (%).
@@ V-C Case Study
In this section, we compare GenBEE and DEGREE based on two examples, which are presented in Figure 4, to illustrate the differences in their predictions. Example 1 presents a case where the DEGREE model incorrectly predicts the argument Site2 in the Binding event triggered by “binding”, but our GenBEE model gives the correct prediction. We infer that our model has incorporated dependency information between arguments using event prompts for enhancing the performance of biomedical event extraction. Example 2 presents a case where the DEGREE model fails to predict the arguments Theme and Site in the Phosphorylation event, which is nested with the Binding event. However, the GenBEE model is able to predict the arguments of these two nested events correctly. We infer that our model has incorporated event structural information using structural prefixes for effectively recognizing the relationships between nested events. As illustrated by the case study, with event prompts and structural prefixes, our proposed GenBEE model is able to perform biomedical event extraction effectively.
@@ VI Conclusion
In this paper, we propose a novel event structure-aware generative model GenBEE, which can effectively tackle complex event structures such as nested events and overlapping events, for biomedical event extraction. The experimental results demonstrate that GenBEE outperforms strong baselines on the MLEE, GE11, and PHEE datasets. In addition, our experiments also show that structure-aware prefixes can effectively serve as a medium to link structural prompts with the representation space of generative models, thereby enhancing the model’s overall performance.